# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 6.1 Reliability of the a posteriori error estimator
$\displaystyle\|\texttt{e}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{\boldsymbol{u}}\|_{0,\Omega}+\|\texttt{e}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}\leq\,C_{\mathrm{rel}}\,\Xi.$
Proof 6.14.
It follows directly from Lemmas 6.9 and 6.11, using the fact that ${\boldsymbol{\varphi}}_{h}$ satisfies the estimate (4.14), and the assumption (6.9).
@@ 6.2 Efficiency of the a posteriori error estimator
(6.10) $C_{\mathrm{eff}}\,\Xi\,\leq\,\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};\Omega}+\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{0,\Omega}+\|\boldsymbol{\rho}-\boldsymbol{\rho}_{h}\|_{0,\Omega}+\|{\boldsymbol{\varphi}}-{\boldsymbol{\varphi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p-p_{h}\|_{0,\Omega}+{\rm h.o.t},$
@@
$\|\boldsymbol{f}+\mathbf{div}\boldsymbol{\sigma}_{h}\|_{0,K}\lesssim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};K}\,{\quad\hbox{and}\quad}\|\boldsymbol{\sigma}_{h}-\boldsymbol{\sigma}^{\tt t}_{h}\|_{0,K}\lesssim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};K}.$
Proof 6.17.
By employing the same arguments as in [17, Theorem 3.2], we can conclude that $\boldsymbol{f}=-\mathbf{div}\,\boldsymbol{\sigma}$, which, together with the symmetry of $\boldsymbol{\sigma}$, implies the desired result. Further details are omitted.
In order to derive the upper bounds for the remaining terms defining the error estimator $\Xi_{s,K}$, we use results from [19], inverse inequalities, and the localisation technique based on element-bubble and edge-bubble functions. The main properties that we will use can be found in [32, Lemmas 4.4-4.7].
@@
$\begin{array}[]{ll}h_{K}\|\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}-\boldsymbol{\nabla}\boldsymbol{u}_{h}+\boldsymbol{\rho}_{h}\|_{0,K}\\ \displaystyle\qquad\lesssim h_{K}\big{(}\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};K}+\|\boldsymbol{\rho}-\boldsymbol{\rho}_{h}\|_{0,K}+\|p-p_{h}\|_{0,K}\big{)}+\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{0,K}.\end{array}$
Proof 6.19.
@@
$\begin{array}[]{ll}h_{K}\|\underline{\mathbf{curl}}(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\|_{0,K}\lesssim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};K}+\|\boldsymbol{\rho}-\boldsymbol{\rho}_{h}\|_{0,K}+\|p-p_{h}\|_{0,K},\\[12.91663pt] h_{E}^{1/2}\|[\![(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\boldsymbol{s}]\!]\|_{0,E}\lesssim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};\omega_{E}}+\|\boldsymbol{\rho}-\boldsymbol{\rho}_{h}\|_{0,\omega_{E}}+\|p-p_{h}\|_{0,\omega_{E}},\end{array}$
@@
Proof 6.21.
[…]
@@
Assume that $\boldsymbol{u}_{\mathrm{D}}$ is piecewise polynomial. Then, for all $E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})$, there holds
$\begin{array}[]{ll}h_{E}^{1/2}\|(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\boldsymbol{s}-\frac{\mathrm{d}\boldsymbol{u}_{\mathrm{D}}}{\mathrm{d}\boldsymbol{s}}\|_{0,E}\lesssim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};K_{E}}+\|\boldsymbol{\rho}-\boldsymbol{\rho}_{h}\|_{0,K_{E}}+\|p-p_{h}\|_{0,K_{E}},\\[12.91663pt] h_{E}^{1/2}\|\boldsymbol{u}_{\mathrm{D}}-\boldsymbol{u}_{h}\|_{0,E}\lesssim h_{K_{E}}\big{(}\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{\mathbf{div};K_{E}}+\|\boldsymbol{\rho}-\boldsymbol{\rho}_{h}\|_{0,K_{E}}+\|p-p_{h}\|_{0,K_{E}}\big{)}+\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{0,K_{E}},\end{array}$
@@
Proof 6.23.
[…]
A direct application of Lemmas 6.16-6.22 yields
(6.11) $\sum_{K\in\mathcal{T}_{h}}\Xi_{s,K}\lesssim\|\texttt{e}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{\boldsymbol{u}}\|_{0,\Omega}+\|\texttt{e}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}.$
Similarly, using the same arguments as in Lemmas 6.16-6.22, along with algebraic manipulations as in Section 5, assuming that $p_{\mathrm{D}}$ is piecewise polynomial, together with the Lipschitz continuity of $\kappa$ (cf. (2.3)), we can bound each of the terms that appear in the estimator $\Xi_{f,K}$ and obtain the following result
(6.12) $\sum_{K\in\mathcal{T}_{h}}\Xi_{f,K}\lesssim\|\texttt{e}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{\boldsymbol{u}}\|_{0,\Omega}+\|\texttt{e}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}.$
@@
For simplicity, we have assumed that $\boldsymbol{u}_{\mathrm{D}}$ and $p_{\mathrm{D}}$ are piecewise polynomial in the derivation of (6.11) and (6.12). However, similar estimates can also be obtained by assuming $\boldsymbol{u}_{\mathrm{D}}$ and $p_{\mathrm{D}}$ are sufficiently smooth (taking, for example, $\boldsymbol{u}_{\mathrm{D}}\in\mathbf{H}^{1}(\Gamma_{\mathrm{D}})$ and $p_{\mathrm{D}}\in\mathrm{H}^{1}(\Gamma_{\mathrm{D}})$, as in Lemmas 6.9-6.11), and proceeding as in [23, Section 6.2]. In such cases, higher-order terms, stemming from errors in the polynomial approximations, would appear in (6.11) and (6.12), accounting for the presence of ${\rm h.o.t}$ in (6.10).
@@
We conclude this section by noting that the a posteriori error estimation analysis developed here can be adapted to the three-dimensional case. In particular, in [31, Th 3.2] and [25, Lemma 4.4], one can find the suitable Helmholtz decompositions for the spaces $\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)$ and $\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)$, respectively.
@@ 7 Numerical results
The computational examples in this section verify the theoretical properties (optimal convergence, conservativity, and robustness) of the proposed schemes. The implementation has been carried out using the FE library FEniCS [1]. The nonlinear systems were solved with Newton–Raphson’s method with a residual tolerance of $10^{-7}$, and the linear systems were solved using the sparse LU factorisation of MUMPS [4].
@@ 7.1 Optimal convergence to smooth solutions and conservativity in 2D
We first consider a simple planar problem setup with manufactured exact solution. We take the unit square domain $\Omega=(0,1)^{2}$, the bottom and left segments represent $\Gamma_{\mathrm{D}}$ and the top and right sides are $\Gamma_{\mathrm{N}}$. We choose the body load $\boldsymbol{f}$, mass source $g$, boundary displacement $\boldsymbol{u}_{\mathrm{D}}$, boundary pressure $p_{\mathrm{D}}$, as well as (not necessarily homogeneous, but standard arguments can be used to extend the theory to the inhomogeneous case) boundary data ${\boldsymbol{\varphi}}\cdot\mbox{\boldmath{$n$}}=\varphi_{\mathrm{N}}$ and $\boldsymbol{\sigma}\mbox{\boldmath{$n$}}=\sigma_{\mathrm{N}}$, such that the exact displacement and fluid pressure are
$\boldsymbol{u}(x,y)=\frac{1}{20}\begin{pmatrix}\cos\bigl{[}\frac{3\pi}{2}(x+y)\bigr{]}\\[4.30554pt] \sin\bigl{[}\frac{3\pi}{2}(x-y)\bigr{]}\end{pmatrix},\quad p(x,y)=\sin(\pi x)\sin(\pi y).$
These exact primary variables are used to construct exact mixed variables of stress, rotation, and discharge flux. We choose the second constitutive relation for the permeability in (1.8) and choose the following arbitrary model parameters (all adimensional) $k_{0}=k_{1}=c_{0}=\alpha=0.1,\quad\lambda=\mu=\mu_{f}=1$. These values indicate a mild permeability variation and it is expected that the nonlinear solver (in this case, Newton–Raphson) converges in only a few iterations. We construct six levels of uniform mesh refinement of the domain, on which we compute approximate solutions and the associated errors for each primal and mixed variable in their natural norms. Convergence rates are calculated as usual:
@@
$k$ DoFs $h$ $e(\boldsymbol{\sigma})$ rate $e(\boldsymbol{u})$ rate $e(\boldsymbol{\rho})$ rate $e({\boldsymbol{\varphi}})$ rate $e(p)$ rate
@@
We also exemplify the momentum and mass conservativity of the formulation. To do so we represent the loss of momentum and mass as
$\texttt{mom}_{h}:=\bigl{\|}\boldsymbol{\mathcal{P}}_{h}[\mathbf{div}(\boldsymbol{\sigma}_{h})+\boldsymbol{f}]\bigr{\|}_{\ell^{\infty}},\qquad\texttt{mass}_{h}:=\biggl{\|}\mathcal{P}_{h}\biggl{[}\bigl{(}c_{0}+\frac{d\alpha^{2}}{d\lambda+2\mu}\bigr{)}p_{h}+\frac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\boldsymbol{\sigma}_{h}+\mathop{\rm div}\nolimits({\boldsymbol{\varphi}}_{h})+g\biggr{]}\biggr{\|}_{\ell^{\infty}},$
where $\mathcal{P}_{h}:\mathrm{L}^{2}(\Omega)\to\mathrm{P}_{k}(\mathcal{T}_{h})$ is the scalar version of $\boldsymbol{\mathcal{P}}_{h}$. They are computed at each refinement level and tabulated in Table 7.2 together with the total error $\texttt{e}:=e(\boldsymbol{\sigma})+e(\boldsymbol{u})+e(\boldsymbol{\rho})+e({\boldsymbol{\varphi}})+e(p)$, and its experimental convergence rate. We report on the nonlinear iteration count as well. The expected optimal convergence of the total error, and the announced local conservativity are confirmed. We also note that, at least for this parameter regime, for all the refinements and polynomial degrees the nonlinear solver takes three iterations to get a residual below the tolerance. In the last column of the same table we report on the efficiency of the global a posteriori error estimator designed in Section 6 $\texttt{eff}(\Xi)=\frac{\texttt{e}}{\Xi}$, which – in this case of smooth solutions – is asymptotically constant (approximately 0.98 for $k=0$ and 1.52 for $k=1$).
DoFs $h$ e rate momh massh iter eff($\Xi$)
@@
Table 7.2: Example 1. Total error, experimental rates of convergence, $\ell^{\infty}$-norm of the projected residual of the momentum and mass balance equations, Newton–Raphson iteration count, and efficiency index of the global a posteriori error estimator. Tabulated results correspond to the two lowest-order polynomial degrees.

### v2

@@ 6.1 Reliability of the a posteriori error estimator
$\displaystyle\|\texttt{e}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{{\bm{u}}}\|_{0,\Omega}+\|\texttt{e}_{\bm{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}\leq\,C_{\mathrm{rel}}\,\Xi.$
Proof.
It follows directly from Lemmas 6.5 and 6.6, using the fact that ${\bm{\varphi}}_{h}$ satisfies the estimate (4.1), and the assumption (6.6). ∎
@@ 6.2 Efficiency of the a posteriori error estimator
$C_{\mathrm{eff}}\,\Xi\,\leq\,\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};\Omega}+\|{\bm{u}}-{\bm{u}}_{h}\|_{0,\Omega}+\|\bm{\rho}-\bm{\rho}_{h}\|_{0,\Omega}+\|{\bm{\varphi}}-{\bm{\varphi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}+\|p-p_{h}\|_{0,\Omega}+{\rm h.o.t},$ (6.7)
@@
$\|\bm{f}+\mathbf{div}\bm{\sigma}_{h}\|_{0,K}\lesssim\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};K}\,{\quad\hbox{and}\quad}\|\bm{\sigma}_{h}-\bm{\sigma}^{\tt t}_{h}\|_{0,K}\lesssim\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};K}.$
Proof.
By employing the same arguments as in [17, Theorem 3.2], we can conclude that $\bm{f}=-\mathbf{div}\,\bm{\sigma}$, which, together with the symmetry of $\bm{\sigma}$, implies the desired result. Further details are omitted. ∎
In order to derive the upper bounds for the remaining terms defining the error estimator $\Xi_{s,K}$, we use results from [19], inverse inequalities, and the localisation technique based on element-bubble and edge-bubble functions. The main properties that we will use can be found in [33, Lemmas 4.4-4.7].
@@
$\begin{array}[]{ll}h_{K}\|\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}-\bm{\nabla}{\bm{u}}_{h}+\bm{\rho}_{h}\|_{0,K}\\ \displaystyle\qquad\lesssim h_{K}\big{(}\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};K}+\|\bm{\rho}-\bm{\rho}_{h}\|_{0,K}+\|p-p_{h}\|_{0,K}\big{)}+\|{\bm{u}}-{\bm{u}}_{h}\|_{0,K}.\end{array}$
Proof.
@@
$\begin{array}[]{ll}h_{K}\|\underline{\mathbf{curl}}(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\|_{0,K}\lesssim\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};K}+\|\bm{\rho}-\bm{\rho}_{h}\|_{0,K}+\|p-p_{h}\|_{0,K},\\[12.91663pt] h_{E}^{1/2}\|[\![(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\bm{s}]\!]\|_{0,E}\lesssim\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};\omega_{E}}+\|\bm{\rho}-\bm{\rho}_{h}\|_{0,\omega_{E}}+\|p-p_{h}\|_{0,\omega_{E}},\end{array}$
@@
Proof.
It suffices to apply [33, Lemma 4.7] with $\bm{\xi}:=\mathcal{C}^{-1}\bm{\sigma}+\dfrac{\alpha}{d\lambda+2\mu}p\mathbf{I}+\bm{\rho}=\bm{\nabla}{\bm{u}}$ and $\bm{\xi}_{h}:=\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h}$. ∎
@@
Assume that ${\bm{u}}_{\mathrm{D}}$ is piecewise polynomial. Then, for all $E\in\mathcal{E}_{h}(\Gamma_{\mathrm{D}})$, there holds
$\begin{array}[]{ll}h_{E}^{1/2}\|(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\bm{s}-\frac{\mathrm{d}{\bm{u}}_{\mathrm{D}}}{\mathrm{d}\bm{s}}\|_{0,E}\lesssim\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};K_{E}}\!+\!\|\bm{\rho}-\bm{\rho}_{h}\|_{0,K_{E}}\!+\!\|p-p_{h}\|_{0,K_{E}},\\[8.61108pt] h_{E}^{1/2}\|{\bm{u}}_{\mathrm{D}}-{\bm{u}}_{h}\|_{0,E}\lesssim h_{K_{E}}\big{(}\|\bm{\sigma}-\bm{\sigma}_{h}\|_{\mathbf{div};K_{E}}+\|\bm{\rho}-\bm{\rho}_{h}\|_{0,K_{E}}+\|p-p_{h}\|_{0,K_{E}}\big{)}+\|{\bm{u}}-{\bm{u}}_{h}\|_{0,K_{E}},\end{array}$
@@
Proof.
The first estimate follows as in [33, Lemma 4.18], defining $\bm{\xi}$ and $\bm{\xi}_{h}$ as in Lemma 6.11. On the other hand, the second estimate follows from an application of the discrete trace inequality (see [33, Lemma 4.6]), using that $\mathcal{C}^{-1}\bm{\sigma}+\dfrac{\alpha}{d\lambda+2\mu}p\mathbf{I}=\bm{\nabla}{\bm{u}}-\bm{\rho}$, and the fact that ${\bm{u}}={\bm{u}}_{\mathrm{D}}$ on $\Gamma_{\mathrm{D}}$. See also [35, Lemma 4.14]. ∎
A direct application of Lemmas 6.9-6.12 yields
$\sum_{K\in\mathcal{T}_{h}}\Xi_{s,K}\lesssim\|\texttt{e}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{{\bm{u}}}\|_{0,\Omega}+\|\texttt{e}_{\bm{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}.$ (6.8)
Similarly, using the same arguments as in Lemmas 6.9-6.12, along with algebraic manipulations as in Section 5, assuming that $p_{\mathrm{D}}$ is piecewise polynomial, together with the Lipschitz continuity of ${\bm{\kappa}}$ (cf. (2.3)), we can bound each of the terms that appear in the estimator $\Xi_{f,K}$ and obtain the following result
$\sum_{K\in\mathcal{T}_{h}}\Xi_{f,K}\lesssim\|\texttt{e}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{{\bm{u}}}\|_{0,\Omega}+\|\texttt{e}_{\bm{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}.$ (6.9)
@@
For simplicity, we have assumed that ${\bm{u}}_{\mathrm{D}}$ and $p_{\mathrm{D}}$ are piecewise polynomial in the derivation of (6.8) and (6.9). However, similar estimates can also be obtained by assuming ${\bm{u}}_{\mathrm{D}}$ and $p_{\mathrm{D}}$ are sufficiently smooth (taking, for example, ${\bm{u}}_{\mathrm{D}}\in\mathbf{H}^{1}(\Gamma_{\mathrm{D}})$ and $p_{\mathrm{D}}\in\mathrm{H}^{1}(\Gamma_{\mathrm{D}})$, as in Lemmas 6.5-6.6), and proceeding as in [23, Section 6.2]. In such cases, higher-order terms, stemming from errors in the polynomial approximations, would appear in (6.8) and (6.9), accounting for the presence of ${\rm h.o.t}$ in (6.7).
@@
We conclude this section by noting that the a posteriori error estimation analysis developed here can be adapted to the three-dimensional case. In particular, in [32, Th 3.2] and [25, Lemma 4.4], one can find the suitable Helmholtz decompositions for the spaces $\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)$ and $\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)$, respectively.
@@ 7 Numerical results
The computational examples in this section verify the theoretical properties (optimal convergence, conservativity, and parameter robustness) of the proposed schemes. The implementation has been carried out using the FE library FEniCS [1]. The nonlinear systems were solved with Newton–Raphson’s method with a residual tolerance of $10^{-7}$, and the linear systems were solved using the sparse LU factorisation of MUMPS [4].
@@ 7.1 Optimal convergence to smooth solutions and conservativity in 2D
We first consider a simple planar problem setup with manufactured exact solution. We take the unit square domain $\Omega=(0,1)^{2}$, the bottom and left segments represent $\Gamma_{\mathrm{D}}$ and the top and right sides are $\Gamma_{\mathrm{N}}$. We choose the body load $\bm{f}$, mass source $g$, boundary displacement ${\bm{u}}_{\mathrm{D}}$, boundary pressure $p_{\mathrm{D}}$, as well as (not necessarily homogeneous, but standard arguments can be used to extend the theory to the inhomogeneous case) boundary data ${\bm{\varphi}}\cdot\mbox{\boldmath{$n$}}=\varphi_{\mathrm{N}}$ and $\bm{\sigma}\mbox{\boldmath{$n$}}=\bm{\sigma}_{\mathrm{N}}$, such that the exact displacement and fluid pressure are
${\bm{u}}(x,y)=\frac{1}{20}\begin{pmatrix}\cos\bigl{[}\frac{3\pi}{2}(x+y)\bigr{]}\\[4.30554pt] \sin\bigl{[}\frac{3\pi}{2}(x-y)\bigr{]}\end{pmatrix},\quad p(x,y)=\sin(\pi x)\sin(\pi y).$
These exact primary variables are used to construct exact mixed variables of stress, rotation, and discharge flux. We choose the second constitutive relation for the permeability in (1.8) and choose the following arbitrary model parameters (all nondimensional) $k_{0}=k_{1}=c_{0}=\alpha=0.1,\quad\lambda=\mu=\mu_{f}=1$. These values indicate a mild permeability variation and it is expected that the nonlinear solver (in this case, Newton–Raphson) converges in only a few iterations. We construct six levels of uniform mesh refinement of the domain, on which we compute approximate solutions and the associated errors for each primal and mixed variable in their natural norms. Convergence rates are calculated as usual:
@@
$k$ DoFs $h$ $e(\bm{\sigma})$ rate $e({\bm{u}})$ rate $e(\bm{\rho})$ rate $e({\bm{\varphi}})$ rate $e(p)$ rate
@@
We also exemplify the equilibrium and mass conservativity of the formulation. To do so we represent the loss of equilibrium and mass as
${\texttt{equ}_{h}}:=\bigl{\|}\bm{\mathcal{P}}_{h}[\mathbf{div}(\bm{\sigma}_{h})+\bm{f}]\bigr{\|}_{\ell^{\infty}},\ \texttt{mass}_{h}:=\biggl{\|}\mathcal{P}_{h}\biggl{[}\bigl{(}c_{0}+\frac{d\alpha^{2}}{d\lambda+2\mu}\bigr{)}p_{h}+\frac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\bm{\sigma}_{h}+\mathop{\rm div}\nolimits({\bm{\varphi}}_{h})+g\biggr{]}\biggr{\|}_{\ell^{\infty}},$
where $\mathcal{P}_{h}:\mathrm{L}^{2}(\Omega)\to\mathrm{P}_{k}(\mathcal{T}_{h})$ is the scalar version of $\bm{\mathcal{P}}_{h}$. They are computed at each refinement level and tabulated in Table 7.2 together with the total error $\texttt{e}:=e(\bm{\sigma})+e({\bm{u}})+e(\bm{\rho})+e({\bm{\varphi}})+e(p)$, and its experimental convergence rate. We report on the nonlinear iteration count as well. The expected optimal convergence of the total error, and the announced local conservativity are confirmed. We also note that, at least for this parameter regime, for all the refinements and polynomial degrees the nonlinear solver takes three iterations to get a residual below the tolerance. In the last column of the same table we report on the efficiency of the global a posteriori error estimator designed in Section 6 $\texttt{eff}(\Xi)=\frac{\texttt{e}}{\Xi}$, which – in this case of smooth solutions – is asymptotically constant (approximately 0.98 for $k=0$ and 1.52 for $k=1$).
DoFs $h$ e rate equh massh iter eff($\Xi$)
@@
Table 7.2: Example 1. Total error, experimental rates of convergence, $\ell^{\infty}$-norm of the projected residual of the equilibrium and mass balance equations, Newton–Raphson iteration count, and efficiency index of the global a posteriori error estimator. Tabulated results correspond to the two lowest-order polynomial degrees.